OpenStudy (anonymous):

Help! How do I simplify this rational expression? :'( 10x/x(2x-1) - 6(x-2)/5(x-2)

OpenStudy (anonymous):

\[10x/x(2x-1) - 6(x-2)/5(x-2)\] \[= (5)10x(x-2)/5x(2x-1)(x-2) - 6x(x-2)(2x-1)/5x(x-2)(2x-1)\] \[= 50x(x-2)-6x(x-2)(2x-1)/5x(2x-1)(x-2)\]

OpenStudy (anonymous):

i don't even know anymore D;

OpenStudy (lgbasallote):

is it written like \[\LARGE \frac{10x}{x(2x-1)} - \frac{6(x-2)}{5(x-2)}???\]

OpenStudy (anonymous):

yup!

OpenStudy (lgbasallote):

cancel terms first

OpenStudy (lgbasallote):

\[\LARGE \frac{10}{2x - 1} - \frac{6}{5}\]

OpenStudy (lgbasallote):

i canceled x in the first term...and (x-2) in the second term

OpenStudy (anonymous):

ohh.. okay

OpenStudy (lgbasallote):

so now..can you operate that?

OpenStudy (anonymous):

wait a second

OpenStudy (lgbasallote):

suree

OpenStudy (anonymous):

ohhh i got it but one question, for the 2nd last step: why is it a positive 6 and not a negative 6?

OpenStudy (anonymous):

ohhh i got it but one question, for the 2nd last step: why is it a positive 6 and not a negative 6? \[= 50/10x-5 - 12x+6/10x-5\] \[= 50-12x+6/10\] I thought it would be a negative..

OpenStudy (lgbasallote):

sorry got logged off...do you still need an explanation?

OpenStudy (anonymous):

yes please :$

OpenStudy (lgbasallote):

so you take the LCD... \[\LARGE \frac{5(10) - 6(2x - 1)}{5(2x-1)}\] did you get this?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

now we simplify...5 x 10 is 50 distribute -6 to (2x - 1) -6 times 2x = -12x -6 times -1 = 6 therefore you get \[\Large \frac{50 - 12x + 6}{10x - 5}\] got it?

OpenStudy (anonymous):

ohh it's -6 not positive six.. okay, got it! thank you!

OpenStudy (lgbasallote):

great!! ^_^